OpenStudy (anonymous):

how do you integrate (s+1)/(4-s^2)^.5 from s=0 to s=2? is the answer supposed to be -2 or pi-2 and do you use u substitution or trig substitution?

OpenStudy (anonymous):

\[\int\limits(s+1)/(4-s ^{2})^{5} ds\] is this the question?

OpenStudy (anonymous):

oops i thought (s+1) is in the denom

OpenStudy (anonymous):

here u have to simply express (s+1) =A(-2s)+ B (-2s is obtained after differentiating (4-s^2) wrt s )

OpenStudy (abb0t):

How is your answer -2 or π/2 for an indefinite integral? o.0

OpenStudy (abb0t):

And I wouldn't suggest partial decomp since it looks like it's rasied to -0.5 power, which means that it's a square root.

OpenStudy (anonymous):

@nitz please don't get confused

OpenStudy (anonymous):

when i got confused???

OpenStudy (anonymous):

ur integral reduces to -1/2 integral of (-2s/(4-s^2)^.5) + integral of (1/(4-s^2)^.5) now first part can be integrated using 4-s^2 =z^2 and for the second part u have standard formulae

OpenStudy (raden):

it similar :